Carbohydrates = 4.1 Calories/gram (includes sugars and starches; insoluble fiber cannot

be digested by humans, and does not provide calories for humans)

Protein = 4.1 Calories/gram

Lipids = 9.5 Calories/gram

1. A one cup serving of cereal has 23 g of total carbohydrates (where one gram of

carbohydrates is insoluble fiber) and two grams of protein. How many calories are in this

serving?
Answer #1

ANSWER.

* Calculation of calories present in the serving.

     Carbohydrate present in the serving = 23 gm ( 22 gm of sugar and starch +

                                                                                 1 gm of insoluble fiber).

          Calories in 1 gm of carbohydrate = 4.1 calories.

          Insoluble fiber not give any calories.

     So

       Calories present in 23 gm of serving = 22 * 4.1

                                                                 = 90.2 calorie.

Protein present in the serving = 2 gm.

              Calorie in 1 gm of protein = 4.1 calorie.

        So

           Calorie present in 2 gm of protein = 2 * 4.1

                                                                  = 8.2 calorie.

        Total calorie present in the serving = 90.2 + 8.2

                                                                = 98.4 calories.